OpenStudy (anonymous):

Find the product: 1/2 times 3/7 how to find the product?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*\frac{ 3 }{ 7 } = \frac{ 3 }{ ? }\] Can you figure it out?

OpenStudy (anonymous):

14 but do i have to simplify the 7 first?

OpenStudy (anonymous):

No, you just multiply all the way across. So your answers should be: \[\frac{ 3 }{ 14 }\] And this is already in the simplest form, so this should be your final answer. I hope this helped :o)

OpenStudy (anonymous):

thanks other question but this one u divide 3/5 divided 2/10 = 3/5 ???

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } / \frac{ 2 }{ 10 }\] Okay, so for this one we need a common denominator which can be 10. So, we're going to multiple 3/2 by 2/2 and 2/10 we'll leave alone. \[\frac{ 3 }{ 5 }*\frac{ 2 }{ 2 }=\frac{ 6 }{ 10 }\] \[\frac{ 6 }{ 10 }\div \frac{ 2 }{ 10 }\] Now we're just going to divide. So 6/2= 3 and 10/10=1 So, \[\frac{ 3 }{ 1 } = 3\] :o)

OpenStudy (anonymous):

oh had to find a common denominator o get now last question also dividing... 3/4 divide 4/5 have to find the write quotient in simplest form. so is it.... 6/5??

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }\div \frac{ 4 }{ 5 }\]\[\frac{ 3 }{ 4 }*\frac{ 5 }{ 4 } =\frac{ 15 }{ 16 }\] I forgot about this. When you divide by a fraction, you take the reciprocal of the second fraction and multiply. :o)

OpenStudy (anonymous):

for dividing fraction theres a little chant: dividing by fractions is as easy as pie, just flip the second and multiply so for the first one (3/5)/ (2/10) u flip 2/10 and get 10/2 then u multiply (3/5) * (10/2) and get 30/10 which is 3

OpenStudy (anonymous):

i like your riddle i rote it down @ bronco101 and for butterfly16 can u remind me what a reciprocal is?

OpenStudy (anonymous):

Of course, a reciprocal is basically the opposite of a fraction. The numbers basically switch. Heres a few examples: \[\frac{ 3 }{ 4 }\rightarrow \frac{ 4 }{ 3 }\]\[\frac{ 12 }{ 20 }\rightarrow \frac{ 20 }{ 12 }\] Do you remember? :o)